\begin{thebibliography}{10}

\bibitem{bakke2022maximal}
Karen Bakke~Haga, David Schrittesser, and Asger Törnquist.
\newblock Maximal almost disjoint families, determinacy, and forcing.
\newblock {\em Journal of Mathematical Logic}, 22(01):2150026, 2022.

\bibitem{ellentuck1974new}
Erik Ellentuck.
\newblock A new proof that analytic sets are {R}amsey.
\newblock {\em The Journal of Symbolic Logic}, 39(1):163--165, 1974.

\bibitem{galvin1973borel}
Fred Galvin and Karel Prikry.
\newblock Borel sets and {R}amsey's theorem.
\newblock {\em The Journal of Symbolic Logic}, 38(2):193--198, 1973.

\bibitem{hruvsak2017ramsey}
Michael Hru{\v{s}}{\'a}k, David Meza-Alc{\'a}ntara, Egbert Th{\"u}mmel, and
  Carlos Uzc{\'a}tegui.
\newblock Ramsey type properties of ideals.
\newblock {\em Annals of Pure and Applied Logic}, 168(11):2022--2049, 2017.

\bibitem{kahane1992operations}
Sylvain Kahane.
\newblock Op{\'e}rations de hausdorff it{\'e}r{\'e}es et r{\'e}unions
  croissantes de compacts.
\newblock {\em Fundamenta Mathematicae}, 141(2):169--194, 1992.

\bibitem{kechris2025theory}
Alexander~S Kechris.
\newblock {\em The theory of countable Borel equivalence relations}.
\newblock Cambridge University Press, 2025.

\bibitem{kechris2006axiom}
Alexander~S Kechris, Robert~M Solovay, and John~R Steel.
\newblock The axiom of determinacy and the prewellordering property.
\newblock In {\em Cabal Seminar 77--79: Proceedings, Caltech-UCLA Logic Seminar
  1977--79}, pages 101--125. Springer, 2006.

\bibitem{mathiasthesis}
Adrian Richard~David Mathias.
\newblock {\em On a generalization of {R}amsey's theorem.}
\newblock PhD thesis, University of Cambridge, 1969.

\bibitem{MATHIAS197759}
Adrian Richard~David Mathias.
\newblock Happy families.
\newblock {\em Annals of Mathematical Logic}, 12(1):59--111, 1977.

\bibitem{schrittesser2019ramsey}
David Schrittesser and Asger T{\"o}rnquist.
\newblock The {R}amsey property implies no mad families.
\newblock {\em Proceedings of the National Academy of Sciences},
  116(38):18883--18887, 2019.

\bibitem{silver}
Jack Silver.
\newblock Every analytic set is {R}amsey.
\newblock {\em The Journal of Symbolic Logic}, 35(1):60--64, 1970.

\bibitem{solovay1970model}
Robert~M. Solovay.
\newblock A model of set-theory in which every set of reals is lebesgue
  measurable.
\newblock {\em Annals of Mathematics}, 92(1):1--56, 1970.

\end{thebibliography}

\documentclass[11pt,letterpaper]{amsart}
\usepackage{amsmath} 
\usepackage{amsthm}
\renewcommand{\mspace}{[\mathbb N]^\infty}
\DeclareMathOperator{\Runif}{\mathsf{Runif}}
\newtheorem{thm}{Theorem}

\newcommand{\hset}{w}

\newcommand{\zfaxiom}{{\sf ZF}}

\title[Ramsey regularity implies no MAD families without uniformization]{Ramsey regularity implies no MAD families without uniformization}

	\author{Jialiang He}
\address{College of Mathematics, Sichuan University, Chengdu, Sichuan, 610064 China;}
\email{jialianghe@scu.edu.cn}

\author{Jintao Luo}
\address{College of Mathematics, Sichuan University, Chengdu, Sichuan, 610064 China;}
\email{jintaoluo@foxmail.com}
\author{Shuguo Zhang}

\address{College of Mathematics, Sichuan University, Chengdu, Sichuan, 610064 China;}
\email{zhangsg@scu.edu.cn}

\begin{document}

	\begin{abstract}
		We show that if the Ramsey property holds (in a class of sets), then there is no MAD family (in this class, provided it satisfies some modest closure properties), proving a conjecture made by A.R.D.\ Mathias in 1977. As the technique we introduce for this proof is useful in a variety of related problems, we take the opportunity to announce 4 theorems, which will be proved in a follow-up paper.
	\end{abstract}
	\maketitle
	
	\section{Introduction}
	
	It is a well-known phenomenon that sets such as the paradoxical decomposition of the sphere, Vitali sets, and other sets constructed with the help of the Axiom of Choice, cannot have a simple description in the sense of definitional complexity. For instance, analytic sets in the sense of the projective hierarchy are regular in the sense of Lebesgue measurable, thereby ruling out the sets mentioned above from appearing at this level of definability. 

Mathias conjectured in 1977 that the regularity property known as the Ramsey property, in a similar manner, rules out the existence of certain irregular sets known as MAD families:

A subset $X$ of $\mspace$, the Polish space of infinite subsets of the natural numbers, has \emph{Ramsey property}, or is \emph{Ramsey regular}, if there exists a set $w \in \mspace$ such that $[w]^\infty \subseteq X$, or $[w]^\infty \cap X = \emptyset$. This regularity property extrapolates to subsets of $\mspace$, the behavior of colorings of finite sets of natural numbers as described by Ramsey's Theorem. 
Every analytic set has the Ramsey property \cite{galvin1973borel,silver,ellentuck1974new}. 

A set $\mathcal{A}\subseteq\mspace$ is called a MAD family if $\mathcal A$ is infinite, for all distinct $A, B \in \mathcal{A}$, $A \cap B$ is finite, and for every $C \in \mspace$ there exists some $A \in \mathcal{A}$ such that $C\cap A$ is infinite.	
MAD families exist by the Axiom of Choice, and
motivated by his extension \cite{mathiasthesis,MATHIAS197759} of results of Solovay \cite{solovay1970model} regarding Lebesgue measurability and the Baire property,
Mathias suspected that they could not exist in the realm of sets which are Ramsey regular, leading to his conjecture \cite[p.~87]{MATHIAS197759}.

To state our main theorem, recall that by a boldface pointclass one means a collection of subsets of Polish spaces which is closed under continuous preimages and containing all clopen sets \cite[p.~118]{kechris2006axiom}.
For our purposes, call a pointclass \emph{good} if it is also closed under intersections and continuous images.

\begin{thm}
If $\Gamma$ is a good pointclass and every set in $\Gamma \cap \mspace$ has the Ramsey property, then there is no MAD family in $\Gamma\cap\mspace$. 
\end{thm}

Mathias' \cite{mathiasthesis,MATHIAS197759} does prove that the conclusion holds for $\Gamma$ equal to the analytic sets.
This is where the matter stood until Schrittesser and Törnquist in 2019 showed that supposing, in addition to `Every set in $\Gamma$ has the Ramsey property', a principle they called $\Runif(\Gamma)$, indeed there are no MAD families in $\Gamma$ \cite{schrittesser2019ramsey}.

Building on their ideas, we remove the spurious assumption $\Runif(\Gamma)$, and are moreover able to give a short and elementary proof of Theorem~1.

\section{Proof of Mathias' conjecture}

We prove Theorem~1 in the following, slightly more general form:
\begin{thm} 	Let $\Gamma$ be a good pointclass.
If every set in $\Gamma$ has the Ramsey property, then there is no Dedekind infinite MAD family in $\Gamma$.

\end{thm}
A set $A$ is \emph{Dedekind infinite} if there exists an injective function  $f: \mathbb N \to A$. Our proof of the above theorem uses only $\zfaxiom$; the Axiom of Choice for countable subsets of $\mspace$ suffices to guarantee that every MAD family is Dedekind infinite.

\begin{proof}
Let $\mathcal A \in \Gamma$ be a Dedekind infinite MAD family, and fix a sequence $\{A_n: n\in \mathbb{N}\}$ of distinct elements in $\mathcal A$.
By replacing each $A_n$ with $A_n\setminus \bigcup_{i<n} A_i$ if necessary, we may assume without loss of generality that the $A_n$ are pairwise disjoint.

For any $x\in\mspace$, we denote by $x(k)$ the $k$-th element of  $x$ under its increasing enumeration.
For $z\in \mspace$, define:
\[\widetilde{z}=\{A_{z(2n)}(z(2n+1)): n\in\mathbb{N}\}.\]

We need two claims:

\textbf{Claim 1}: 
For every $\hset\in\mspace$,  there exist $y\in [\hset]^\infty$	and $A\in \mathcal A$
such that $ \widetilde{y}\subseteq A$.

\begin{proof} Since $\mathcal A$ is a MAD family and $\widetilde{\hset}$ is an infinite subset of $\mathbb{N}$,  there exists some $A\in \mathcal A$ such that $\widetilde{\hset}\cap A$ is infinite. 
	By the definition of $\widetilde{\hset}$, we have $A\not= A_n$ for all $n\in \mathbb{N}$.
	Enumerate $\widetilde{\hset}\cap A$ as:
	\[ \{A_{\hset(2n_k)}(\hset(2n_k+1)): k\in\mathbb{N} \}\]
	Now, define: 
	\[
	y=\{\hset(2n_k):k\in\mathbb{N}\}
	\cup\{\hset(2n_k+1):k\in\mathbb{N}\}.
	\]  
	Then, we have \[\widetilde{y}=\{A_{y(2n)}(y(2n+1)):n\in\mathbb{N}\}=\widetilde{\hset}\cap A\subseteq A,\]
	as required.
\end{proof}

\textbf{Claim 2}: 
For every $\hset\in \mspace$,  there exist $y\in [\hset]^\infty$	and $A\in \mathcal A$
such that both $ A\cap\widetilde{y}$ and $ \widetilde{y}\setminus A$ are infinite.		

\begin{proof} 
	By the previous claim, we can assume that $\widetilde{\hset} \subseteq A$, for some $A \in \mathcal A$.
	As $A\not= A_n$ for all $n\in \mathbb{N}$, $A\cap A_n$ is finite for all $n\in\mathbb{N}$.

	We construct $y \subseteq \hset$ by giving a recursive definition of its enumeration:
	Assuming we have already defined $y(j)$ for $j<i$ and that $i$ is divisible by $4$, we define the next four elements, $y(i), \dots, y(i+3)$. With $\hset_e= \{\hset(2n) : n \in \mathbb{N}\}$,
	\begin{itemize}
		\item Find $y(i) \in \hset_e$, requiring $y(i) > y(i-1)$ in case $i>0$,
		\item Find $y(i+1) \in \hset$ large enough so that $A_{y(i)}\big(y(i+1)\big) \notin A$; this is possible as $A_{y(i)} \cap A$ is finite;
		\item Find $y(i+2) \in \hset_e$ such that $y(i+2) > y(i+1)$
		and let $y(i+3)$ be the next element of $\hset$  after $y(i+2)$, which ensures that  
		$A_{y(i+2)}\big(y(i+3)\big) \in A$.
	\end{itemize}
	Clearly $\widetilde{y}=\big\{A_{y(2i)}\big(y(2i+1)\big):i\in\mathbb{N}\big\}$ has the   required property.
\end{proof}	

Let us define \[P=\{z\in \mspace: \exists A\in \mathcal A (\widetilde{z}\subseteq A)\}.\] 

Since $\Gamma$ is a good pointclass, $P\in \Gamma$. 
We now show $P$ is not Ramsey regular. 
Otherwise, one of the following cases obtains:

\textbf{Case 1}: Suppose that there exists $\hset \in \mspace$ such that $[\hset]^\infty \cap P = \emptyset$.
By Claim 1, there are $y \in [\hset]^\infty$ and $A \in \mathcal A$ with $\widetilde{y} \subseteq A$, i.e., $y \in P$.
This contradicts the assumption that $[\hset]^\infty \cap P=\emptyset$.

\textbf{Case 2}:  Suppose there exists $\hset \in \mspace$ such that $[\hset]^\infty \subseteq P$.
By Claim 2, there exist $y \in [\hset]^\infty$ and $A \in \mathcal A$ such that both $A\cap \widetilde{y}$ and $\widetilde{y}\setminus A$ are infinite. We show that $y \notin P$: Since $\mathcal A$ is maximal, there exists $B \in \mathcal A$ with $ (\widetilde{y}\setminus A)\cap B$ infinite.
Because $(\widetilde{y}\setminus A)\cap B \neq \emptyset$, it follows that $A \neq B$.
Note that an element of $P$ can have an infinite intersection with at most one element of $\mathcal A$. Since both $A$ and $B$ have infinite intersection with $\widetilde{y}$, it follows that $y \notin P$.
This contradicts the assumption that $[\hset]^\infty \subseteq P$.

Both cases lead to a contradiction, completing the proof.
\end{proof}

\section{Further results}

We conclude by announcing several further results which can be shown by the same method. These will be proved in a follow-up paper. 

\begin{thm}	
Suppose $\Gamma$ is a good pointclass.
If every set in $\Gamma$ has the Ramsey property, then 

\begin{enumerate}
	\item There is no Vitali set (and hence no Hamel bases) in $\Gamma$.
	\item There is no maximal independent family in $\Gamma$.	
	\item There is no  Dedekind infinite $\mathcal{ED}$-MAD family  in $\Gamma$, and also no Dedekind infinite $\mathcal{ED}_{fin}$-MAD family (by almost the same proof).	
	\item For each $\alpha<\omega_1$, there is no  Dedekind infinite $\mathbf{Fin^\alpha} $-MAD family in the pointclass $\Gamma$.	
\end{enumerate}
\end{thm}

Item (4) above was already shown in \cite{bakke2022maximal} under the additional assumption $\Runif(\Gamma)$, where the reader will also find the definition of the notion of $\mathcal I$-MAD family for an arbitrary ideal $\mathcal I$ on $\mathbb N$. 
The ideal $\mathbf{Fin^\alpha}$ was defined by Kahane \cite{kahane1992operations}; see also \cite[p.~100]{kechris2025theory}.
The ideals $\mathcal{ED}$ and $\mathcal{ED}_{fin}$ are defined, e.g., in \cite{hruvsak2017ramsey}.
Note that as a corollary of the above theorem, in particular there are no analytic 
$\mathcal{ED}$-MAD or $\mathcal{ED}_{fin}$-MAD families.

	\section*{Acknowledgments}

The authors thank David Schrittesser and Hang Zhang for many valuable comments and discussions. The authors gratefully acknowledges supported by  the  National Natural Science Foundation of China (11801386).


	\bibliographystyle{plain}
\bibliography{mybibtex}
\end{document}